Three charges are sitting at the corners of an equilateral triangle. If ๐‘ž1 = ๐‘ž3 = 10 ๐œ‡๐ถ , ๐‘ž2 = 5 ๐œ‡๐ถ and the length of each side is 20 cm, find the force experienced by charge ๐‘ž2 in unit-vector notation and magnitude-angle notation.


1
Expert's answer
2021-03-09T15:29:29-0500
"F_x=0\ N\\F=F_y=2F'\sin{60}
\\F=F_y=2(9\cdot10^9)\frac{(10\cdot10^{-6})(5\cdot10^{-6})}{0.2^2}\sin{60}=19.5\ N"

"\vec{F}=19.5\ N \vec{j}\\
\vec{F}=19.5\ N \text{ 90}\degree"